You own Huntington Beach Boats, and just received a purchase for a $1.5 million yacht. Unfortunately, you have no money in your business account and need to borrow to build the boat! It will cost you $800,000 to build and you will be able to pay back the money in a year. Three banks you have visited are willing to give you the money with the boat as collateral if the buyer does not pay. CSUSB Credit Union is offering you an $800,000 loan at 12% simple interest. Huntington Beach Bank has offered you an $800,0000 loan at 11.8% compounded monthly. Finally, Bank of Laguna has offered you an 11% loan compounded continuously, for the same amount. Whose loan should you go with and what is the EAR of this loan? What is the APR of each loan?
Answer #1

APR:
CSUSB =12%
Huntington Beach Bank=11.8%
Bank of Laguna=11%

EAR:
CSUSB=12%
Huntington Beach Bank=(1+11.8%/12)^12-1=12.46%
Bank of Laguna=e^(11%)-1=11.63%

Go ahead with Bank of Laguna as it has lowest EAR